### v2

# Studying Differential Mental Health Expressions in India

Abstract
Psychosocial stressors and the symptomatology of mental disorders vary across cultures. However, current understandings of mental health expressions on social media are predominantly derived from studies in WEIRD (Western, Educated, Industrialized, Rich, and Democratic) contexts. In this paper, we analyze mental health posts on Reddit made by individuals in India, to identify variations in online depression language specific to the Indian context compared to users from the Rest of the World (ROW). Unlike in Western samples, we observe that mental health discussions in India additionally express sadness, use negation, are present-focused, and are related to work and achievement. Illness is uniquely correlated to India, indicating the association between depression and physical health in Indian patients. Two clinical psychologists validated the findings from social media posts and found 95% of the top-20 topics associated with mental health discussions as prevalent in Indians. Significant linguistic variations in online mental health-related language in India compared to ROW, emphasize the importance of developing precision-targeted interventions that are culturally appropriate.
Studying Differential Mental Health Expressions in India
**footnotetext: These authors served as co-first authors, alphabetically ordered by last name.
1 Introduction
Over 197 million individuals in India are diagnosed with mental health disorders Sagar et al. (2020), a disproportionate majority of whom do not receive treatment Singh (2018). The treatment gap for mental health disorders goes up to 95% in India, which is the highest across Asian countries and more severe compared to the gap of 78% in the United States (US) Murthy (2017); Naveed et al. (2020). Stigma, limited accessibility to mental healthcare, and a lack of awareness of mental health disorders, cumulatively contribute to diagnostic barriers for mental health care Meshvara (2002); Lahariya (2018); Krendl and Pescosolido (2020).
Cognitive impairments such as self-rumination and negative evaluation owing to depression can be measured from language. For instance, increasing use of the first-person singular pronouns reflects self-rumination Spasojević and Alloy (2001); Holtzman et al. (2017), and phrases expressing self-deprecation indicate negative self-evaluation Kovacs and Beck (1978). Social media platforms are increasingly used for mental health-related conversations and seeking support in India Akbar et al. (2020). Language in social media posts can be harnessed to measure cultural differences in the manifestation of depression  Chancellor and De Choudhury (2020); Guntuku et al. (2017); Eichstaedt et al. (2018). Reddit, in particular, provides unique affordances, such as the choice to be anonymous, for individuals to openly discuss their experiences without any character limit and seek support in dedicated threads De Choudhury and De (2014); Boettcher (2021). Language-based AI models for screening depression can facilitate timely, non-intrusive targeted assessments, including support and treatment, especially in under-resourced contexts such as India, alleviating the challenges associated with traditional assessment methods  World Health Organization (2021).
While it is long known that depression manifests differently across cultures Manson (1995), but, how is still an unanswered research question. Lacking awareness of cultural differences in the manifestation of depression leads to misdiagnosis even in clinical settings Bailey et al. (2009); Lewis-Fernandez et al. (2005). More recently, Rai et al. (2024) demonstrated differences in the language of White and Black individuals with depression, raising questions about the generalizability of widely acknowledged cognitive impairments across populations. This study is the first to focus on depression expressions in social media language posted by users in India. More importantly, we collaborate with clinical psychologists in India to validate trends and themes obtained from social media posts, which bolsters this computational study. A related previous study by De Choudhury et al. (2017) grouped India with other developing nations - for example, South Africa.Another study on understanding Indian population work by Pendse et al. (2019), examined help-seeking patterns rather than depression markers, to inform social media platform design.
Understanding how the language of people with depression varies across cultures calls for interpretable features such as n-grams and Linguistic Inquiry and Word Count (LIWC) Boyd et al. (2022) relevant to other stakeholders such as psychologists Aguirre et al. (2021); Loveys et al. (2018); Burkhardt et al. (2022); Rai et al. (2024). Using a blend of cross-disciplinary language features and machine learning models, we address the overarching question, if and how the mental health expressions of Indian users on social media are different from the rest of the world by answering the following:
  • How do linguistic expressions in Reddit posts of individuals experiencing mental health challenges in India differ from individuals in the rest of the world?
  • How well do data-driven insights on mental health expressions align with the experience of clinical psychologists in India?
  • Do the linguistic expressions of depression specific to India differ enough to reliably distinguish individuals from MH-India compared to those from MH-RoW?
This paper bridges two critical gaps from previous literature. First, the paper identifies mental health expressions specific to India, the world’s most populous country, by mining Reddit threads. Second, it engages with clinical psychologists practicing in India to validate the empirical findings, providing a culturally informed assessment of cross-country comparisons of mental health expressions.
2 Data
Reddit offers a platform for individuals to share their mental health journey and seek support anonymously, thereby making it a rich source to understand the symptomatology and sequelae of mental health De Choudhury and De (2014); Park et al. (2018); Gaur et al. (2018); Lokala et al. (2022). Other more widely used social media platforms such as Mx Takatak, Moj, Tiktok in India lack APIs for data collection oos . Previously, Reddit posts have been used for identifying shifts to suicidal ideation De Choudhury et al. (2016), depression symptoms Gaur et al. (2018); Liu et al. (2023), and the mental health expressions of immigrants Mittal et al. (2023b), among others.
Figure 1: The count of users for each country in the Rest of World control group (log scale). Demonstrates that the large majority of users in the ROW group are geolocated to Western countries. The "Others" Category contains countries with less than 10 users, including Belgium (9), Italy (9), Mexico (6), Malaysia (5), Romania (4), Croatia (4), UAE (2), South Africa (2), China (2), Spain (2), Greece (2), Denmark (1), Finland (1), Iceland (1), Japan (1), South Korea (1), Poland (1), Russia (1), Singapore (1), Thailand (1), Turkey (1) and Vietnam (1).
2.1 Subreddits: Mental Health vs Control
We extracted $3,195,310$ posts and comments from mental health-related subreddits (See Appendix A for list of subreddits) using the PushShift API Baumgartner et al. (2020). The largest portion of users (36.1%) were members of r/depression. The subreddits for the Control group were queried across all remaining subreddits external to the mental health subreddits.
Most posts and comments in the dataset were posted between 2019 and 2020. As a preprocessing step, we removed deleted usernames and null messages. Assuming that users posting in India-specific subreddits are likely to be Indians, we identified India-focused subreddits (See Appendix A) and then grouped together the users who posted in these subreddits. After this step and filtering users who posted at least 500 words (excluding comments), we grouped users into four groups:
  • MH-India (4185 users): “Individuals geolocated in India and posting in Mental Health Subreddits”,
  • MH-ROW (5588 users): “Individuals geolocated outside India (i.e., Rest of World) and posting in Mental Health Subreddits”,
  • Control-India (2622 users): “Individuals geolocated in India and not posting in mental health subreddits” and,
  • Control-ROW (5594 users): “Individuals geolocated outside India and not posting in mental health subreddits”.
The first group (MH-India) is our group of interest; the remaining are controls. We specifically chose posts as the scope of this study was to obtain expressions of experiences with mental health challenges rather than interactions with others’ mental health challenges.
2.2 Geolocation - India vs ROW
We used the geolocation inference approach Harrigian (2018) as a second layer of verification for user location. The geolocation model is a location estimation model that utilizes word usage, the frequency distribution of subreddit submissions, and the temporal posting habits of each user to determine their location. Specifically, we use the pre-trained GLOBAL inference model [^1] https://github.com/kharrigian/smgeo/tree/master#models to geolocate users in our dataset. We removed any users not geolocated to their group based on subreddit classification. For example, users in MH-India who are not geolocated to India and users in MH-RoW who are geolocated to India were removed. This functioned as a two-step verification to ensure that users in MH-India were from India. Ultimately, 1200 users out of the initial 4185 users were left in the MH-India group, and 930 users out of 2622 were left in the Control-India group. Most users in the ROW group were geolocated to the US (See Fig 1), affirming the dominance of West-centric data on Reddit.
We evaluated the quality of geolocation by manually verifying the self-disclosed location for randomly sampled 100 users. We found that the model’s estimate of the individual’s country matched the self-disclosed location, even though the state or city estimate was not always accurate.
2.3 Matching Control groups with users in MH-India
Age and gender are well-known confounders in behavioral health studies Schwartz et al. (2013). We first estimated the age and gender of every user in our dataset using a machine-learning approach described in Appendix B. We then matched the users from our group of interest, i.e., MH-India, with those in control groups (MH-ROW, and Control-ROW) on these two covariates. Owing to the smaller sample size, matching was not performed for the Control-India group. However, the age distribution across the four groups was fairly similar before matching, with the average age being 25 for the MH-India, Control-India, and Control-ROW groups and 24 for the MH-ROW group.
Ideally, the focus and control group samples should have indiscernible covariates. However, exact matching Rosenbaum (2020) is difficult to achieve without dropping a large set of samples. Coarsened Exact Matching (CEM) Iacus et al. (2009) is a softer version of Exact Matching, which stretches the matching criteria wide enough to avoid dropping samples that are similar but not an exact match. We implement CEM using MatchIt package Stuart et al. (2011) in R and set the distance to ‘Mahalanobis’ for one-to-one matching. The quality of matching was evaluated using Standard Mean Differences and Kolmogorov-Smirnov Statistics (See Appendix C). The mean age was 24.7 (sd= 3.41). The mean gender score was -0.97 (sd= 0.93), where a higher positive score indicates female. Table 1 shows the total number of posts and users in each of the four groups after CEM.
Group # Distinct Users # Posts
MH-India 1200 50928
Control-India 930 69957
MH-ROW 1200 54666
Control-ROW 1200 122654
Total 4530 298205
Table 1: Number of users and posts in each of the four groups of our dataset.
3 Methods
3.1 Language Features
We extracted a diverse set of interpretable language features widely used in Psychology literature to understand the unique markers of mental health among Indians. This includes n-grams, psychosocial word categories from LIWC, and topics (word clusters derived using LDA).
  1. 1.
    We extracted 1-3 grams from posts and created a normalized bag-of-words representation for each user. We filtered out 1-3 grams having point-wise mutual information (PMI)$\leq$5. N-grams reveal cultural idioms of distress Desai and Chaturvedi (2017) such as tension Weaver (2017) that are unique to a culture.
  2. 2.
    LIWC-22 is a dictionary comprising 102 word categories based on psycho-social states (e.g., Cognition, Social Processes, Affect, etc.) to measure mental states from the language. These word categories in LIWC are counted for each user, and the count is normalized by the total number of 1-grams for each user, thereby representing each user as a vector of 102 normalized psychosocial categories.
  3. 3.
    We used Latent Dirichlet Allocation (LDA) Blei et al. (2003) to extract latent topics in users’ timeline data. While recent neural topic modeling methods such as BERTopic Grootendorst (2022) show superior predictive accuracy, LDA has been shown to provide qualitatively robust topics Dixon et al. (2022).
    We generated three sets of topics by setting the number of topics $=[200,500,2000]$. We evaluated the topics’ quality using Topic Uniqueness (TU) Nan et al. (2019). TU represents the number of times a set of keywords is repeated across topics; a higher TU corresponds to a rarely repeated word, indicating that topics are diverse, which is favorable. Additionally, three co-authors independently reviewed the quality of topics. We set the number of topics to 2000 based on the automated (See Table A3) and manual evaluation. High granularity in topics is desired as it reveals subtle variations in language across cultures.
3.2 Correlation Analysis
To measure the association between language and the groups (i.e., MH-India vs MH-ROW), we performed ordinary least squares regressions with the feature sets. We calculated Pearson $r$ to measure the association of each feature to each group in a one-vs-all setting. This is important in supporting an understanding of how each feature set correlates with the presence of mental health challenges in India. p-values were corrected using Benjamini-Hochberg correction for multiple hypothesis testing. 102 word categories for LIWC, 2000 for LDA topics, and 23,344 1-3 grams were considered for p-value correction.
3.3 Predictive Model
To examine whether the linguistic expressions of the MH-India group sufficiently differ to discriminate it from other groups, we trained ‘one vs rest’ logistic regression models in a 10-fold cross-validation setting Rifkin and Klautau (2004). More sophisticated methods (such as XGBoost) could potentially provide higher performance, but the focus of the study is not to achieve state-of-the-art performance for group prediction but to test if sufficient discriminating evidence exists across groups to warrant culturally aligned depression models. We report the Area Under the Receiver Operating Curves (AUC) for each feature for the MH-India and MH-ROW groups.
4 Results
[…]
4.1 Mental Health Expression: India vs ROW
4.1.1 N-grams
Sixty-one 1-3 grams were significantly ($p<0.05$) correlated with the MH-India group, and 156 were correlated with the MH-ROW group. Figure 2 illustrates the top 25 1-3 grams arranged in decreasing order of Pearson $r$ for both groups. First-person singular pronouns (‘i am’, ‘i’), and mentions of depression (‘depressed’, ‘depression’), are prevalent in both groups. Help-seeking phrases (‘suggest’, ‘advice’, ‘talk’, ‘please help’) are prevalent in MH-India, whereas MH-ROW discusses feelings, use clinical language (‘diagnosed’, ‘medication’) and express negative emotions (‘anxiety’, ‘bad’, ‘worse’) associated with depression. Interestingly, the MH-India group simultaneously discusses the feelings of loneliness and social relationships with parents and friends. Academic-related stress (‘college’, ‘exam’, ‘study’) is exclusively seen in discussions of the MH-India group. This is particularly interesting, considering users in both groups were matched for age, yet discussions around student-life challenges are prevalent only in MH-India.
Overall, the discussion in MH-India subreddits is centered around seeking help, whereas negative feelings are more commonly discussed in the MH-ROW group.
4.1.2 LIWC
Fifty-two LIWC categories were significantly associated ($p<0.05$) with the MH-India group, whereas sixty categories were found to be correlated with the MH-ROW group. We provide the Top 20 LIWC categories for both groups in Table 2. While negative tone, anxiety, and personal pronouns are correlated with depression in both groups, sadness, an additional affect is uniquely seen in the MH-Indian group. Depression-related discourse in MH-ROW group is past-focused, which is a known marker of depression whereas it is present-focused in India. Social behavioral attributes (communication, politness), work, and achievement are correlated with depression in the MH-India group but not in MH-ROW. Somatic symptoms/illness (pain, sick) are seen in the MH-India group, whereas substance abuse/addiction is correlated with the MH-ROW group.
MH-India MH-ROW
Theme Top Words r Theme Top Words r
Negative/ Suicidal thoughts life, parents, family, hate, die 0.280 Negative Feelings feel, myself, feeling, depression, anymore 0.358
Depression mentions anymore, depression, tired, depressed, everyday 0.229 Negative Feelings went, didn, crying, mad, stayed 0.340
Negative Feelings feel, myself, feeling, depression, anymore 0.227 Depression mentions anymore, depression, tired, depressed, everyday 0.338
Friends & Relationships friends, talk, social, anxiety, alone 0.226 Negative Feelings feeling, body, feels, heart, scared 0.322
Friends & Relationships love, heart, loved, beautiful, happiness 0.207 Symptoms sick, woke, stomach, switched, asleep 0.318
Table 3: Top Words and Correlation Coefficient (r) by Theme for MH-India and MH-ROW
4.1.3 LDA Topics
Of $2000$ topics, $109$ were found to be significant ($p<0.05$) for the MH-India group and $216$ for MH-ROW group. The top topics and the corresponding Pearson r, are provided in Table LABEL:tab:topics_output. The most prevalent topics in the MH-India group discussed negative/suicidal thoughts (life, parents, family, hate, die), academic and job stressors ((college, exam, study, university, engineering), job, degree, college, school, career) and family and relationships (love, heart, loved, beautiful, happiness). In contrast, the prevalent themes in MH-ROW are negative feelings and emotions ((feel, myself, feeling, depression, anymore);(went, didn, crying, mad, stayed); (feeling, body, feels, heart, scared)). Topics such as mental disorders (anxiety, depression, mental, medication, disorder), goodbyes (sister, ma, papa, clutching, plane) and anger (hate, angry, rant, respect, ugly) are common in both sets.
4.2 Validation
We validated the significantly correlated topics for our group of interest (MH-India) by showing the top words to two clinical psychologists with significant practical experience with seeing patients and mental healthcare in India. Specifically, we asked the following question:
To what extent the open vocabulary topics having a significant correlation with the MH-India group are prevalent in Indian patients? - A Likert scale of 0-5 is provided where $5$ indicates ‘Highly Prevalant’ and ‘0’ indicates ‘Not observed at all’.
Prevalence While independently labeling topics for prevalence, the clinical psychologists agreed with each other $81.49\%$ of the time. Of the top 20 topics significantly associated with the MH-India group, 95% were ranked either extremely or somewhat prevalent (4 or 5 on a scale of 1 - 5) in India by at least one of the two clinical psychologists, and 80% were ranked as prevalent (a score of 4 or 5) by both evaluators. Of the 109 topics significantly associated with the MH-India group, 56% were annotated as prevalent by at least one evaluator.
4.3 Is the language in MH-India different from control groups?
We used logistic regression model with each feature set (i.e., n-grams, LIWC and LDA topics) to predict membership in MH-India vs MH-ROW. High AuC scores (See Table 4) demonstrate that users’ language in the MH-India group significantly differs from those in the control groups, including MH-ROW. All language feature groups (i.e., n-grams, LIWC, and LDA topics) have fairly high AUC, indicating differences in depression expression at 1-3 gm level, in psycho-social categories as well as in latent topics of discussion.
1-3 grams LIWC LDA Topics
MH-India 0.853 0.776 0.758
MH-ROW 0.881 0.818 0.811
Table 4: AUCs for Logistic regression one vs. rest models predicting group membership.
5 Discussion
Our work reveals significant differences in the language markers of mental health expressions of Indian users compared to those from outside India. The association with politeness coupled with discussions around family, work and achievement indicates that the users in the MH-India group tend to associate mental health with their ability and social relationships as opposed to more mentions of swear words and feelings in the ROW group. Academic and family pressures are unique to Indian users, possibly due to the collectivist nature of Indian society Chadda and Deb (2013). The matching performed across control groups rules out the majority young demographic in social media data as the possible reason. The association with present focus words in MH-India contrasts with the widespread belief of self-focussed rumination when suffering from mental illness Park et al. (2017). MH-India group also tends to reason (‘causation’ words - how, because, why) more in their language in contrast to expressing feelings in MH-ROW group. Communication (phone, call, message, post/tweet/meme, sms/texting, chat) is also exclusive to the MH-India group. We speculate high reliance on social media platforms for mental health support and privacy potentially due to stigma associated with mental health in face-to-face conversations Shidhaye and Kermode (2013).
Our work reaffirms cultural differences such as substance use in the MH-ROW population vs somatic complaints in the MH-India population. Unlike in Western cultures where substance use is a barrier to mental health support, it is familial concerns and the fear of stigma in MH-India Biswas et al. (2016). We speculate based on these findings that culturally-aware mental health intervention techniques should seek to bridge the gap to treatment by addressing hypochondriacal ideas and familial embarrassment in particular.
Only $56\%$ of 109 topics correlated with the MH-India group were labeled as prevalent in Indian patients by clinical psychologists. We speculate that some of the topics not labeled as "prevalent" are unseen or emerging themes. Of Top-20 topics in MH-India, "not prevalent" topics revolve around Video Games/Online Content, Grooming/Physical Appearance, and Programming, indicating the influence of digital content, growing isolation, and low self-esteem amongst the undiagnosed young population. These topics could be underrecognized concerns. The second group of topics includes Environmental Impact of Energy Sources, Humorous content and reactions, among others. Previous research has suggested that people (particularly young people) are increasingly climate anxious and that humor on social media is often used to cope with mental health challenges Schneider (2018); Sanson (2022). Furthermore, certain pop culture references may be crucial to understanding the narrative - as one clinical psychologist pointed out, the top word "Singh" in one of the topics may correspond to the suicide of late Bollywood star Sushant Singh Rajput, a significant event that potentially catalyzed a range of important conversations surrounding mental health in India Akbar et al. (2020).
The growing treatment gap for mental disorders is a major concern in Indian society. The economic loss from mental health conditions between 2012-2030 is estimated at USD 1.03 trillion [^2] United Nations: https://www.who.int/india/health-topics/mental-health. Automated systems that could diagnose and support mental well-being can potentially alleviate the lack of resources, but they would only be useful when designed considering the cultural sensitivities and norms of society. This study establishes significant linguistic variations in the mental health-related language in social media posts by Indians compared to individuals from the rest of the world. It thus becomes imperative to test the computational depression detection models for cultural invariance before their deployment in clinical settings Aguirre et al. (2021); Rai et al. (2023) and indicate the need for socio-culturally aware mental health models to prevent misdiagnosis.
6 Background
Depression and anxiety disorders are the most imminent mental health challenges, with the highest contribution to Indian Disability Adjusted Life Years  Sagar et al. (2020). Fear of embarrassing family members is a primary barrier to mental health recovery in India whereas, for example, substance abuse is the major hurdle in America  Biswas et al. (2016). $71\%$ of Indians exhibited stigma when answering questions about mental health Foundation (2018). Relatedly, somatic symptoms, hypochondriasis, anxiety, and agitation are more commonly seen in Indian patients compared to psychological symptoms  Gada (1982). While the extent of mental health stigma and treatment (un)availability is often studied, it remains unknown how individuals suffering from mental disorders in India express and seek support on social media.
There is accumulating evidence that suggests language markers of depression vary with demographics such as race Rai et al. (2024); Aguirre and Dredze (2021), immigrant status Mittal et al. (2023a), and geographic location De Choudhury et al. (2017). For instance, Indian and South Africa-based users are less candid in their posts and less likely to exhibit negative emotions in comparison to their Western counterparts De Choudhury et al. (2017). Another study studied the language of Asian users including Indian, Malaysian, and Filipino on Mental Health Support Forum, Talklife Pendse et al. (2019) and found that Indians discuss “wanting or needing friends” more than other countries. .
Limitations
The text-based geolocation of individuals in this study could potentially label Indians who later moved to other countries as Indians residing in India. Further, the Reddit user sample does not represent the general population, as evidenced by the mostly English language data in our India samples, although India has over 100 languages. In particular, we note that the majority of users were geolocated to Karnataka (a southern state in India) and that the age (ranging between 12 and 48) distributions are not necessarily representative. Our work shows the significant cultural themes observed in Indian society. However, Reddit posts represent a small population of India. While this analysis provides correlational insight into the data, it does not offer causal claims.
Ethical Considerations
Our university’s Institutional Review Board deemed this study exempt due to the public nature of all data. While Reddit data is public, it may contain users’ personal information, including city and town. We limited our analysis to country and state-level geolocation information to reduce the possibility of personally identifying individuals. Gender was predicted using a continuous scale, with extremes indicating masculinity and feminity. We exercised caution while presenting linguistic patterns and examples not to reveal any individual’s timeline quotes. Members of our team have not viewed or worked with individual-level granular data. When done ethically with respect to user anonymity and privacy, we believe this line of research could assist in understanding diverse individuals’ mental health challenges and developing personalized interventions that improve the well-being and mental health of under-resourced communities Proferes et al. (2021).
Acknowledgment
We acknowledge the limited use of OpenAI’s ChatGPT in the writing and refining this manuscript.
Appendix A Subreddits Used to Extract the Raw Data
Mental Health Subreddits:
The mental health subreddit was obtained from prior works Sharma and De Choudhury (2018); Saha et al. (2020). These include: r/Anxiety, r/bipolar, r/BipolarReddit, r/depression, r/sad, r/SuicideWatch, r/addiction, r/opiates, r/ForeverAlone, r/BPD, r/selfharm, r/StopSelfHarm, r/OpiatesRecovery, r/Sadness, r/schizophrenia, r/AdultSelfHarm
Control Subreddits:
All subreddits excluding Mental health subreddits.
India focused Subreddits:
r/india, r/mumbai, r/tamil, r/Hindi, r/Kerala, r/Urdu, r/delhi, r/pune, r/hyderabad, r/bangalore, r/kolkata, r/telugu, r/marathi, r/AskIndia, r/sanskrit, r/Kochi, r/Rajasthan, r/pali, r/Chandigarh, r/Chennai, r/karnataka, r/Bhopal, r/Coimbatore, r/kannada, r/TamilNadu, r/Trivandrum, r/gujarat, r/punjabi, r/Bengali, r/kolhapur, r/Vijaywada, r/Dehradun, r/sahitya, r/Uttarakhand, r/ahmedabad, r/bharat, r/nagpur, r/Agra, r/assam, r/Indore, r/surat, r/navimumbai, r/Goa, r/sikkim, r/lucknow, r/Bareilly, r/nashik, r/Allahabad, r/Durgapur, r/Jamshedpur, r/Asansol, r/indianews, r/IndianGaming, r/IndiaSpeaks, r/indiameme, r/dankinindia, r/indiasocial
Appendix B Age and Gender
We applied an open-source age and gender predictive lexica Sap et al. (2014) to obtain continuous values of age and gender. This lexicon was built over a set of over 70,000 users from social media and blogs and predicted age with a Pearson r of 0.86 and gender with an accuracy of 0.91 and has been applied reliably on Reddit data in prior studies Zirikly et al. (2019). We used the probabilities from this model to denote the gender attribute of users in our data and did not consider gender as a binary category.
To validate the machine-generated predictions of gender and age for users within the Reddit dataset, we looked for posts containing self-disclosures of gender and age per user. Examples of this include “(23F)” for a user who self identifies as a 23 year old female. We were able to identify 5,844 posts across 706 unique users (See Table A1 for distribution) who employed some form of gender self-identification, allowing us to measure the accuracy of provided gender predictions. Using this subset, the model’s gender prediction holds at 91.89% (See Table A2 for groupwise performance).
Group Name Gender Age
MH India 224 427
MH ROW 195 378
Non-MH India 45 111
Non-MH ROW 242 388
Table A1: Distribution of Users who self-disclosed gender and age by Group
Group Name Accuracy MAE (in yrs)
MH India 87.67% 3.12
MH ROW 86.69% 4.65
Non-MH India 94.33% 4.38
Non-MH ROW 96.12% 6.48
Table A2: Model performance for predicting gender and age. MAE stands for Mean Absolute Error and is reported in years.
Figure 3: Differences in Covariates before and after CEM for groups “Control-ROW” and “MH-ROW”
Appendix C Coarsened Exact Matching
A control group is considered balanced with the treatment group if the difference is close to zero. The focus group MH-India has 1200 users and a total of 1200 users were matched from the groups MH-ROW and Control-ROW. Matching was not performed for Control-India as the number of samples is 930 and we did not want to drop any samples.
Appendix D Topic Uniqueness Scores
Table A3 describes the TU scores generated by each of the tested combinations of number of topics and grouping of Posts and Comments. This was used to choose Posts Only, Number of Topics = 2000 for our analysis.
Message Type Number of Topics Value
Both Messages 200 0.7205
Both Messages 500 0.6194
Both Messages 2000 0.5676
Posts Only 200 0.7675
Posts Only 500 0.7162
Posts Only 2000 0.7694
Table A3: Topic Uniqueness Scores for different numbers of topics and message types.
Appendix E Communication with Clinical Psychologists
Table A3 shows the email text used for communication with each clinical psychologist.
The goal of this project is to study the manifestation of mental illness in Indians. As a part of this project, we have identified a set of 100 Topics/ Themes that Indians Users were found to commonly discuss on Reddit, a social media platform. We have labeled these topics as per our understanding and we now need your help in interpreting these topics from your perspective. The objective is to essentially identify Topic or theme of discussion in the context of mental illness in India,
    How often a theme is observed in an Indian patient suffering from a mental illness?
These identified topics are available in this Google sheet. Please read the below steps carefully:
    Peruse the top words given in Column-A. These are the top 10 common words comprising a single topic.
    In Column E, select the degree of prevalence of this topic amongst Indian patients. The options are Highly prevalent, somewhat prevalent, unsure, rarely observed, and Not observed at all.
    You may add your comments in Column -F
Table A4: Email communication with clinical psychologists who performed an informed review of the topics and ChatGPT-generated labels in this study.